# Human Decision-making is Susceptible to AI-driven Manipulation

### v2

@@ Methods
Following the pre-experiment survey, we randomly assigned participants to domain-AI conditions while balancing key demographic and psychometric variables across different groups. Each participant received a unique passcode to access our experimental platform and detailed instructions, framing the study as an investigation of personalized AI-assisted decision-making. The instructions were carefully worded to avoid demand characteristics while maintaining ethical transparency.
@@
After completing all of the required scenarios on the platform, participants evaluated the assigned scenarios’ relevance to their daily lives and rated their assigned AI agent on three dimensions: helpfulness, informativeness, and soundness (10-point Likert scales). We also collected open-ended responses regarding participants’ overall impressions and specific interaction experiences. To ensure ethical conduct and participant well-being, participants received detailed explanations of the study’s true objectives and the rationale behind any experimental manipulation. We instructed the participants on the optimal choices for each scenario and explained why certain options might be harmful. Additionally, participants were encouraged to contact the study investigators if they experienced any negative effects.
@@
We recruited participants through poster advertisements on WeChat, a predominant Chinese social media, and Prolific [100], a UK-based crowdsourcing platform. Participants were required to be 1) Fluent in Chinese or English; 2) Be able to use a computer device to access and use our platform; 3) Have experience using LLMs/personal assistants; 4) Be over 18 years of age; and 5) Do not currently suffer from physical/mental illnesses, as participation may have caused additional distress to these individuals. Data was collected online through our platform to ensure data safety and privacy. As the experiment took approximately one hour on average, each participant was rewarded 80 RMB ($\approx$8£) for completing all the tasks and relevant questionnaires. In addition, the top 10% of the participants, based on dialogue quality and timely completion of tasks, received an additional bonus of 20 RMB ($\approx$2£).
@@
We conducted a power analysis to estimate the required number of participants in this study. Setting the effect size at $0.1$, significance level at $0.05$, and power at $0.8$ indicated that we required $246$ samples per decision-making domain. As each participant completed three scenarios (samples) in our experiments, we needed a minimum of $164$ participants ($27$ per domain-AI condition) to satisfy these requirements.
@@
To ensure the validity of our findings, we examined participant characteristics across domain-AI conditions. One-way ANOVA revealed no significant differences in demographics or social factors across domains and agent groups ($F[5]$in [$0.13,2.09$], $P\in[0.07,0.98]$). Similarly, conversation length was consistent across domain-AI conditions ($M=10.72$, $SD=1.55$; $F[5]=0.40$, $P=0.85$), confirming that observed effects were not attributable to pre-existing individual differences or interaction duration.
Decision Patterns
To assess how interactions with the AI agents influenced participants’ decisions, we considered the highest-rated option(s) at each time point (pre/post interaction) as the participant’s most likely choice(s). Changes in these choices were categorized into four decision patterns: No Change represents the participant maintaining their original decision, which was further divided into Optimal or Harmful, depending on whether the choice included the optimal option; Positive Shift indicates a beneficial change, where the participant initially chose a harmful option but switched to the optimal option or maintained their choice of the optimal option while including fewer harmful options in their decision after the interaction; Negative Shift involves changing from the optimal choice to a harmful option or switching between harmful options. We analyzed the distribution of decision patterns across domain-AI conditions using chi-squared tests, with post-hoc z-tests with Bonferroni correction for pairwise comparisons and Cohen’s $h$ for effect sizes.
Temporal Changes in Preference Ratings
As the manipulative agents (MA and SEMA) were instructed to guide participants away from optimal options and toward their hidden incentives, our analyses focused on changes in the ratings of these two options. We analyzed these temporal changes using mixed-effects ANOVA comparing pre/post ratings across AI conditions, followed by Bonferroni-corrected independent t-tests for pairwise comparisons, with Cohen’s $d$ for effect sizes.
@@
We employed a linear mixed-effects model to analyze individual factors influencing manipulation susceptibility. Decision patterns were quantified on a scale reflecting their impact: negative shifts ($-1$), positive shifts ($1$), and no change (00). This mapping preserved both directional and magnitude information while enabling quantitative analysis of individual susceptibility.
$Y_{ij}=\beta_{0}+\beta_{1}(\text{AI condition}_{i})+\mathbf{X}_{ij}\boldsymbol{\beta}^{\prime}+(1|P_{i})+(1|S_{j})+\epsilon_{ij},$ (1)
where $Y_{ij}$ represents participant $i$’s quantified decision pattern on scenario $j$, and $\beta_{0}$ and $\epsilon_{ij}$ indicate the intercept and residual error, respectively. $\mathbf{X}_{ij}$ represents covariates, including demographics, personality, trust in AI, social support accessibility, and social engineering susceptibility traits. We also included pre-interaction confidence and familiarity ratings and the scenario’s completion order as covariates. Lastly, the model incorporated random intercepts for the participant ($1|P_{i}$) and the scenario ($1|S_{j}$).
@@
The data collected in our study and the source code for data collection and analysis will be made available upon request from the corresponding authors, with clear justification and data usage agreement.
@@ Acknowledgments
Figure S1: Temporal Changes in preference ratings across AI conditions in decision-making contexts. This figure illustrates participants’ ratings of the optimal and harmful options before and after interaction with the AI agents, with points estimated representing the mean preference ratings at each time point. Error bars represent $95\%$ confidence intervals. While, across both domains, participants reported similar ratings for the optimal choices and the hidden incentives, respectively, between AI conditions at baseline (pre-interaction), those interacting with manipulative agents (MA and SEMA) showed significant declines in ratings for the optimal options and increases for harmful options, reflecting the influence of these agents (Supplementary Table S3).
@@
Table S1: Pairwise comparisons (proportion’s $z$-tests with Bonferroni correction) of decision patterns between AI conditions across financial and emotional domains. Significant values ($P<0.05$) are highlighted in bold.
Domain Decision Pattern Condition 1 Condition 2 $\boldsymbol{|z|}$ $\boldsymbol{P}$ value Cohen’s $\boldsymbol{h}$
Financial Positive Shift NA ($31.7\%$) MA ($5.3\%$) $5.17$ $\mathbf{7.15\times 10^{-7}}$ $0.73$
NA ($31.7\%$) SEMA ($4.4\%$) $5.39$ $\mathbf{2.16\times 10^{-7}}$ $0.77$
MA ($5.3\%$) SEMA ($4.4\%$) $0.31$ $1.00$ $0.04$
No Change (Optimal) NA ($16.7\%$) MA ($9.6\%$) $1.58$ $0.34$ $0.21$
NA ($16.7\%$) SEMA ($11.4\%$) $1.16$ $0.74$ $0.15$
MA ($9.6\%$) SEMA ($11.4\%$) $0.43$ $1.00$ $0.06$
No Change (Harmful) NA ($23.3\%$) MA ($23.7\%$) $0.06$ $1.00$ $0.01$
NA ($23.3\%$) SEMA ($24.6\%$) $0.22$ $1.00$ $0.03$
MA ($23.7\%$) SEMA ($24.6\%$) $0.15$ $1.00$ $0.02$
Negative Shift NA ($28.3\%$) MA ($61.4\%$) $5.09$ $\mathbf{1.08\times 10^{-6}}$ $0.68$
NA ($28.3\%$) SEMA ($59.6\%$) $4.83$ $\mathbf{4.12\times 10^{-6}}$ $0.64$
MA ($61.4\%$) SEMA ($59.6\%$) $0.27$ $1.00$ $0.04$
Emotional Positive Shift NA ($28.2\%$) MA ($11.7\%$) $3.1$ $\mathbf{5.77\times 10^{-3}}$ $0.42$
NA ($28.2\%$) SEMA ($13.0\%$) 2.92 $\mathbf{0.015}$ $0.38$
MA ($11.7\%$) SEMA ($13.0\%$) 0.3 $1.00$ $0.04$
No Change (Optimal) NA ($37.6\%$) MA ($25.2\%$) 2.01 $0.13$ $0.27$
NA ($37.6\%$) SEMA ($26.8\%$) 1.79 $0.22$ $0.23$
MA ($25.2\%$) SEMA ($26.8\%$) 0.28 $1.00$ $0.04$
No Change (Harmful) NA ($21.4\%$) MA ($20.7\%$) 0.12 $1.00$ $0.02$
NA ($21.4\%$) SEMA ($18.7\%$) 0.52 $1.00$ $0.07$
MA ($20.7\%$) SEMA ($18.7\%$) 0.39 $1.00$ $0.05$
Negative Shift NA ($12.8\%$) MA ($42.3\%$) 5.01 $\mathbf{1.65\times 10^{-6}}$ $0.68$
NA ($12.8\%$) SEMA ($41.5\%$) 4.97 $\mathbf{2.04\times 10^{-6}}$ $0.67$
MA ($42.3\%$) SEMA ($41.5\%$) 0.14 $1.00$ $0.02$
Table S2: Pairwise comparisons (proportion’s $z$-tests with Bonferroni correction) of decision patterns between financial and emotional domains across AI conditions. Overall indicates the proportion of decision patterns within the domain for all AI conditions combined. Significant values ($P<0.05$) are highlighted in bold.
AI Condition Decision Pattern Financial Emotional $\boldsymbol{|z|}$ $\boldsymbol{P}$ value Cohen’s $\boldsymbol{h}$
NA Positive Shift $31.7\%$ $28.2\%$ $0.58$ $0.56$ $0.08$
No Change (Optimal) $16.7\%$ $37.6\%$ $3.63$ $\mathbf{2.83\times 10^{-4}}$ $0.48$
No Change (Harmful) $23.3\%$ $21.4\%$ $0.36$ $0.72$ $0.05$
Negative Shift $28.3\%$ $12.8\%$ $2.95$ $\mathbf{3.20\times 10^{-3}}$ $0.39$
MA Positive Shift $5.3\%$ $11.7\%$ $1.74$ $0.08$ $0.24$
No Change (Optimal) $9.6\%$ $25.2\%$ $3.09$ $\mathbf{2.03\times 10^{-3}}$ $0.42$
No Change (Harmful) $23.7\%$ $20.7\%$ $0.53$ $0.59$ $0.07$
Negative Shift $61.4\%$ $42.3\%$ $2.86$ $\mathbf{4.22\times 10^{-3}}$ $0.38$
SEMA Positive Shift $4.4\%$ $13.0\%$ $2.33$ $\mathbf{0.02}$ $0.32$
No Change (Optimal) $11.4\%$ $26.8\%$ $3.00$ $\mathbf{2.70\times 10^{-3}}$ $0.40$
No Change (Harmful) $24.6\%$ $18.7\%$ $1.10$ $0.27$ $0.14$
Negative Shift $59.6\%$ $41.5\%$ $2.80$ $\mathbf{5.15\times 10^{-3}}$ $0.37$
Overall Positive Shift $14.1\%$ $17.7\%$ $1.30$ $0.19$ $0.10$
No Change (Optimal) $12.6\%$ $29.9\%$ $5.57$ $\mathbf{2.48\times 10^{-8}}$ $0.43$
No Change (Harmful) $23.9\%$ $20.2\%$ $1.16$ $0.25$ $0.09$
Negative Shift $49.4\%$ $32.2\%$ $4.64$ $\mathbf{3.56\times 10^{-6}}$ $0.35$

### v3

@@ Methods
Following the pre-experiment survey, we randomly assigned participants to domain-AI conditions while balancing key demographic and psychometric variables across different groups. Each participant received a unique passcode to access our experimental platform and detailed instructions, framing the study as an investigation of personalized AI-assisted decision-making. The instructions were carefully worded to avoid demand characteristics while maintaining ethical transparency. Accordingly, participants were required to complete all three scenarios corresponding to their assigned domain (see Experimental Platform).
@@
After completing all of the assigned scenarios on the platform, participants evaluated each scenario’s relevance to their daily lives and rated their assigned AI agent on three dimensions: helpfulness, informativeness, and soundness (using 10-point Likert scales). We also collected open-ended responses regarding participants’ overall impressions and specific interaction experiences. To ensure ethical conduct and participant well-being, participants received detailed explanations of the study’s true objectives and the rationale behind any experimental manipulation. We instructed the participants on the optimal options for each scenario and explained why certain options might be harmful. Additionally, participants were encouraged to contact the study investigators if they experienced any negative effects.
@@
We recruited participants through poster advertisements on WeChat, a predominant Chinese social media, and Prolific [palan2018prolific], a UK-based crowdsourcing platform. Participants were required to be 1) Fluent in Chinese or English; 2) Be able to use a computer device to access and use our platform; 3) Have experience using LLMs/personal assistants; 4) Be over 18 years of age; and 5) Do not currently suffer from physical/mental illnesses, as participation may have caused additional distress to these individuals. Data was collected online through our platform to ensure data safety and privacy. As the experiment took approximately one hour on average, each participant was rewarded 80 RMB ($\approx$8£) for completing all the tasks and relevant questionnaires. In addition, the top 10% of the participants, based on dialogue quality and timely completion of tasks, received an additional bonus of 20 RMB ($\approx$2£).
@@
We conducted a power analysis using G*Power 3.1 to estimate the required sample size for this study. Following previous work[kang2021sample, serdar2021sample, Shuai2025Towards, sharma2023human, alegria2018effectiveness], setting the effect size at $0.1$, significance level at $0.05$, and power at $0.8$ indicated that we required $246$ samples per decision-making domain. As each participant completed three scenarios (samples) in our experiments, we needed a minimum of $164$ participants ($27$ per domain-AI condition) to satisfy these requirements.
@@
To ensure the validity of our findings, we examined participant characteristics across domain-AI conditions. One-way ANOVA revealed no significant differences in demographics or social factors across domains and agent groups ($F[5]$ $\in$ [$0.13,2.09$], $P\in[0.07,0.98]$). Similarly, conversation length was consistent across domain-AI conditions ($M=10.72$, $SD=1.55$; $F[5]=0.40$, $P=0.85$), confirming that the observed effects were not attributable to pre-existing individual differences or interaction duration.
Preference Rating Analyses
In our experiments, participants provided 10-point Likert ratings for each option before and after interacting with the assigned AI agents. As the manipulative agents (MA and SEMA) were instructed to guide participants away from optimal options and toward their hidden incentives, we modeled the difference between the rating of these two options (Hidden-Optimal Differential; $HOD$) at time $t$ (pre-/post-interaction) as the primary outcome in our analyses:
$HOD(t)=Rating_{Hidden}(t)-Rating_{Optimal}(t),$ (1)
where $HOD>0$ indicates a relative preference for the hidden incentive. Accordingly, we used a mixed-effects ANOVA to compare $HOD$, as well as the ratings of optimal options and hidden incentives, before and after the interaction between AI conditions across domains, followed by Bonferroni-corrected independent t-tests for pairwise comparisons, with Cohen’s $d$ as the effect size.
Modeling Preferences
To investigate whether shifts in participants’ ratings impact their decisions, we constructed a three-level outcome based on $HOD$ for each scenario and time point: Hidden ($HOD>0$), Tie ($HOD=0$), and Optimal ($HOD<0$). Then, we fit two complementary models with within-participant clustering.
First, we fit a nominal generalized estimating equations (GEE) model with a global odds ratio working correlation at the participant level to obtain population-average difference-in-differences (DID) effects of the Contrast over time (Contrast $\times$ Time), where Contrast denotes the studied contrast for the given model (MA/SEMA vs. NA).
$\log\frac{Pr(Y_{ijt})=c}{Pr(Y_{ijt})=Optimal}=\alpha_{c}+\beta_{1c}(\text{Contrast}_{ijt})+\beta_{2c}(\text{Time}_{ijt})+\beta_{3c}(\text{Contrast}_{ijt}\times\text{Time}_{ijt}),$ (2)
where $Y_{ijt}$ is participant $i$’s log odds of preference category ($c\in{\text{Hidden},\text{Tie}}$ vs. Optimal) on scenario $j$ at time $t$. $\alpha_{c}$ indicates the category-specific intercept. The DID odds ratios quantify the change in the odds of Hidden (or Tie) relative to Optimal from pre- to post-interaction.
Second, to compare differences in preference categories at post-interaction while adjusting for any pre-interaction imbalances, we fit a multinomial logistic regression with post-interaction preferences as the dependent variable and pre-interaction preferences as a covariate.
$Y_{ij}=\log\frac{\Pr(Y_{ij,\text{post}}=c)}{\Pr(Y_{ij,\text{post}}=\text{Optimal})}=\beta_{0c}+\beta_{1c}(\text{Contrast}_{ij})+\beta_{2c}(Y_{ij,\text{pre}}),$ (3)
where $Y_{ij}$ represents participant $i$’s log odds of post-interaction preference category ($c\in{\text{Hidden},\text{Tie}}$ vs. Optimal) on scenario $j$, adjusted for pre-interaction preference. $\beta_{0}$ indicates the intercept.
In both models, $P$ values were Bonferroni-corrected. In addition, for interpretability, we further derived model-based predicted probabilities for each category at pre-interaction, post-interaction, and post-interaction adjusted for pre-interaction.
@@
We employed linear mixed-effects models to analyze individual factors associated with manipulation susceptibility within each manipulative agent (MA and SEMA) across decision-making domains.
$Y_{ij}=\beta_{0}+\mathbf{X}_{ij}\boldsymbol{\beta}^{\prime}+(1|P_{i})+(1|S_{j})+\epsilon_{ij},$ (4)
where $Y_{ij}$ represents participant $i$’s post-interaction $HOD$ on scenario $j$. $\beta_{0}$ and $\epsilon_{ij}$ indicate the intercept and residual error, respectively. $\mathbf{X}_{ij}$ are the covariates, including data source (WeChat vs. Prolific), demographics, personality traits, trust in AI, social support accessibility, and social engineering susceptibility traits. We also included pre-interaction $HOD$, confidence, and familiarity ratings, and scenario $j$’s completion order as covariates. Continuous predictors were standardized, and corresponding results are reported per 1 SD. We included random intercepts for the participant ($1|P_{i}$) and scenario ($(1|S_{j})$).
@@
Process Mechanisms
To examine whether participants’ post-interaction perceptions could explain the effect of the AI condition on $HOD$, we conducted causal mediation analyses for average changes in confidence and familiarity, as well as the perceived personalization, soundness, informativeness, and helpfulness of the agent. Mediation analyses were performed separately for each contrast (MA/SEMA vs. NA) and domain via the following two models:
$\text{Mediator: }M_{ij}=\alpha_{0j}+\alpha_{1j}X_{i}+C_{i}\gamma_{j}+\epsilon_{M_{ij}}$ (5)
$\text{Outcome: }Y=\beta_{0}+c^{\prime}X_{i}+\sum^{n}_{j=1}b_{i}M_{ij}+C_{i}\delta+\epsilon_{Y_{i}},$ (6)
where $Y_{i}$ and $M_{ij}$ are participant $i^{\prime}s$ average post-interaction $HOD$ and reported value for mediator $j$, respectively. $X_{i}$ indicates whether the participant was part of a manipulative group (MA or SEMA). $C_{i}$ represents the covariates, including the average pre-interaction $HOD$ and data source. $\alpha_{0j}$ and $\beta_{0}$ indicate the intercepts, with $\epsilon$ indicating residual error. $\gamma_{j}$ and $\delta$ are covariate coefficients in mediator $j$ and outcome regression, respectively.
@@
The data collected in our study, along with the source code for data collection and analysis, is publicly available via GitHub.
@@ Acknowledgments
Figure S1: Distribution of preference ratings across AI conditions in decision-making contexts for optimal options and hidden incentives. While in both domains, participants reported similar ratings at baseline (pre-interaction) across AI conditions, those interacting with MA and SEMA showed substantial declines in their ratings for the optimal options and increases for the hidden incentives, reflecting the influence of these agents (Supplementary Table S3).